### v2

# toltxlabel

Abstract
Longitudinal modified treatment policies (LMTP) have been recently developed as a novel method to define and estimate causal parameters that depend on the natural value of treatment. LMTPs represent an important advancement in causal inference for longitudinal studies as they allow the non-parametric definition and estimation of the joint effect of multiple categorical, ordinal, or continuous treatments measured at several time points. We extend the LMTP methodology to problems in which the outcome is a time-to-event variable subject to a competing event that precludes observation of the event of interest. We present identification results and non-parametric locally efficient estimators that use flexible data-adaptive regression techniques to alleviate model misspecification bias, while retaining important asymptotic properties such as $\sqrt{n}$-consistency. We present an application to the estimation of the effect of the time-to-intubation on acute kidney injury amongst COVID-19 hospitalized patients, where death by other causes is taken to be the competing event.
1 Introduction
In survival analysis, a competing event is one whose realization precludes the occurrence of the event of interest. As an example, consider a study on the effect of mechanical ventilation of hospitalized COVID-19 patients on the onset of acute kidney injury, where the death of a patient by extraneous causes may be considered a competing event. Commonly used methods for analyzing time-to-event outcomes under competing events include estimating the cause-specific hazard function (Prentice et al., 1978), the sub-distribution function (also known as cumulative incidence; Fine & Gray, 1999), or treating the competing event as a censoring event. Early techniques for the estimation of these parameters were developed in the context of the Cox proportional hazards models (Cox, 1972; Andersen & Gill, 1982). More recent methods for analyzing survival data include joint modeling (Henderson et al., 2000) and a multitude of other parametric and semi-parametric methods (e.g., Zeger & Liang, 1992).
While these methods have allowed much progress in applied research, they have two important limitations: (i) they lack a formal framework that clarifies the conditions required for a causal interpretation of the estimates, and (ii) they typically rely upon parametric modeling assumptions that can induce non-negligible bias in the effect estimates, even when the assumptions required for causal identification hold. Recent developments in causal inference and semi-parametric estimation have yielded important progress in addressing these significant limitations. For example, when the treatment is binary, Young et al. (2020) present a study of the definition and identification of several causal effect parameters in the presence of competing risks, and Benkeser et al. (2018) and Rytgaard & van der Laan (2021) describe semi-parametric–efficient estimators that leverage the flexibility of data-adaptive regression procedures in order to mitigate model misspecification bias while retaining asymptotic properties critical to reliable statistical inference (e.g., $\sqrt{n}$-consistency). Among other insights, Young et al. (2020) clarify that cumulative incidence effects may be interpreted as a total effect of exposure operating through pathways that include the competing event, whereas treating the competing event as a censoring event quantifies direct effects operating independent of the competing event. Furthermore, they clarify that identification when treating the competing event as a censoring event entails considering interventions that eliminate the competing event, and therefore require no unmeasured confounding of the competing event–outcome relation. This assumption is not required to identify the cumulative incidence effect, which is the focus of this work. The reader interested in further issues surrounding the definition and usefulness of these effects in different applications is encouraged to read the original research article (Young et al., 2020).
Under a counterfactual causal framework, defining causal effects requires considering outcomes under hypothetical interventions on the treatment of interest. In the case of binary treatments, causal effects are often defined as differences between the expected outcomes under hypothetical interventions that assign the treatment to all units or to none — that is, the average treatment effect. Causal effects defined through such deterministic interventions are limited and or lack a basis for scientific interpretation in many applications, including problems where the treatment of interest is numerical rather than categorical, where the treatment is multivariate, where the treatment is a time-to-event-variable (such as time-to-intubation as in our illustrative application), or when treatment assignment is deterministic conditional on potential confounders (cf. positivity/overlap assumptions).
As a solution to the definition of relevant causal effects for these problems, recent work has focused on stochastic interventions and modified treatment policies (Stock, 1989; Robins et al., 2004; Díaz & van der Laan, 2012; Haneuse & Rotnitzky, 2013; Richardson & Robins, 2013; Young et al., 2014). Stochastic interventions inquire what would have happened in a hypothetical world where the post-intervention treatment is a random draw from a user-specified distribution conditional on covariates. Examples of stochastic interventions include interventional mediation effects (VanderWeele et al., 2014; Díaz et al., 2020; Hejazi et al., 2022c), as well as incremental propensity score interventions whose identification does not require the positivity assumption (Kennedy, 2019; Wen et al., 2021). Modified treatment policies (MTPs) generalize stochastic interventions and allow the post-intervention treatment to also depend on the natural value of treatment (for an overview of such interventions, see Young et al., 2014). MTPs allow the definition and estimation of parameters defined in terms of natural research questions, such as inquiring what the incidence of wheezing among children with asthma would have been had the concentration of PM2.5 particles been 5% lower than they actually were, or what the mortality rate due to opioid overdose would have been had naloxone access laws been implemented a year earlier than they were actually implemented (Rudolph et al., 2021).
Identifying assumptions for causal effects of longitudinal modified treatment policies using the extended g-formula were first articulated by Richardson & Robins (2013) and Young et al. (2014). Recently, Díaz et al. (2021) developed sequential regression and sequentially doubly robust estimators for longitudinal modified treatment policies (LMTP). These estimators are generalizations of estimators developed by Díaz & van der Laan (2018) for MTPs in the single time-point case, and of estimators developed by Luedtke et al. (2017) and Rotnitzky et al. (2017) for static interventions in the multiple time-point case. Here, we extend the LMTP methodology for the estimation of the cumulative incidence under competing events, illustrating how LMTPs can be used to solve problems whose solution has remained elusive in causal inference, such as non-parametric definition and estimation of effects for continuous, multi-valued, time-valued, or multivariate exposures in the context of time-varying exposures. In our motivating example we illustrate estimation of the effect of time-to-intubation on time to acute kidney injury with death by other causes acting as a competing risk.
Our estimation strategy is rooted in recent developments allowing the estimation of causal effects using flexible regression techniques from machine learning (van der Laan & Rose, 2011, 2018; Chernozhukov et al., 2017). Central to this theory is the notion of von-Mises expansion (e.g., von Mises, 1947; van der Vaart, 1998; Robins et al., 2009), which together with sample-splitting techniques (Bickel, 1982; Klaassen, 1987; Zheng & van der Laan, 2011) allows the use of flexible regression methods for the estimation of nuisance parameters while still allowing the construction of efficient and $\sqrt{n}$-consistent estimators.
2 Defining LMTP effects under competing risks
Assume that at each time point $t=1,\ldots\tau$, we measure $X_{t}=(D_{t},Y_{t},L_{t},A_{t},C_{t})$ on each of $i=1,\ldots,n$ study participants, where $L_{t}$ denotes time-varying covariates, $A_{t}$ denotes a vector of treatments at time $t$, $C_{t}$ denotes an indicator of loss-to-follow-up (equal to one if the unit remains on study at time $t+1$ and equal to zero otherwise), $D_{t}$ denotes an indicator of the competing event occurring on or before time $t$, and $Y_{t}$ denotes an indicator of the event of interest not having occurred on or by time $t$. Baseline covariates are included in $L_{1}$. At the end of the study, we measure the outcome of interest $Y_{\tau+1}$. We let $D_{1}=0$ and $Y_{1}=1$, meaning that participants have not experienced the event of interest nor the competing event at the beginning of the study. We let $R_{t}=\mathds{1}\{D_{t}=0,Y_{t}=1\}$ denote an indicator that neither the event of interest nor the competing event have occurred by time $t$. Assume monotone loss-to-follow-up so that $C_{t}=0$ implies $C_{k}=0$ for all $k>t$, in which case all data become degenerate for $k>t$. In this work, we are interested in a scenario where occurrence of the competing event precludes occurrence of the event of interest. For example, in a study looking to assess the effect of hypertension medications on the likelihood of a stroke, death for reasons other than a stroke is a competing event. In this notation, we have $Y_{t}=0$ if $Y_{t-1}=0$, $D_{t}=1$ if $D_{t-1}=1$, and $Y_{t}=0$ if $D_{t}=1$, since the competing event precludes occurrence of the event of interest. Let $X_{i}=(X_{1,i},\ldots,X_{\tau,i})$ denote the observed data for subject $i$ across all time points $t=1,\ldots\tau$. We let $\mathsf{P}f=\int f(x)\,\mathrm{d}\mathsf{P}(x)$ for a given function $f(x)$. We use $\mathsf{P}_{n}$ to denote the empirical distribution of $X_{1},\ldots\,X_{n}$, and assume $\mathsf{P}$ is an element of the nonparametric statistical model defined as all continuous densities on $X$ with respect to a dominating measure $\nu$. We let $\mathsf{E}$ denote the expectation with respect to $\mathsf{P}$, i.e., $\mathsf{E}\{f(X)\}=\int f(x)\,\mathrm{d}\mathsf{P}(x)$. We also let $||f||^{2}$ denote the $L_{2}(\mathsf{P})$ norm $\int f^{2}(x)\,\mathrm{d}\mathsf{P}(x)$. We use $\bar{W}_{t}=(W_{1},\ldots,W_{t})$ to denote the past history of a variable $W$, use $\underaccent{\bar}{W}_{t}=(W_{t},\ldots,W_{\tau})$ to denote the future of a variable, and use $H_{t}=(\bar{D}_{t},\bar{Y}_{t},\bar{L}_{t},\bar{A}_{t-1})$ to denote the history of the time-varying covariates and the treatment process until just before time $t$. $\bar{W}$ denotes the entire history $(W_{1},\ldots,W_{\tau})$ of a variable. By convention, variables with an index $t\leq 0$ are defined as the null set, expectations conditioning on a null set are marginal, products of the type $\prod_{t=k}^{k-1}b_{t}$ and $\prod_{t=0}^{0}b_{t}$ are equal to one, and sums of the type $\sum_{t=k}^{k-1}b_{t}$ and $\sum_{t=0}^{0}b_{t}$ are equal to zero. For vectors $u$ and $v$, $u\leq v$ denotes point-wise inequalities.
We formalize the definition of the causal effects using a non-parametric structural equation model (Pearl, 2000). Specifically, for each time point $t$, we assume the existence of deterministic functions $f_{D,t}$, $f_{Y,t}$, $f_{L,t}$, $f_{A,t}$, $f_{C,t}$, such that
$D_{t}$ $=f_{D,t}(C_{t-1},A_{t-1},H_{t-1},U_{D,t})$
$Y_{t}$ $=f_{Y,t}(D_{t},C_{t-1},A_{t-1},H_{t-1},U_{Y,t})$
$L_{t}$ $=f_{L,t}(Y_{t},D_{t},C_{t-1},A_{t-1},H_{t-1},U_{Y,t})$
$A_{t}$ $=f_{A,t}(H_{t},U_{A,t})$
$C_{t}$ $=f_{C,t}(A_{t},H_{t},U_{C,t}).$
Here $U=(U_{L,t},U_{A,t},U_{C,t},U_{D,t},U_{Y,t}:t\in\{1,\ldots,\tau+1\})$ is a vector of exogenous variables. Independence assumptions on these errors necessary for identification will be clarified in §2.1. Without loss of generality, we assume that variables which are undefined are equal to zero (e.g., $L_{t}=0$ if $C_{k}=1$ for any $k<t$).
We are concerned with the definition and estimation of the causal effect of an intervention on the treatment process $\bar{A}$ on the event indicator $Y_{\tau+1}$ in a hypothetical world where there is no loss to follow-up (i.e., $\bar{C}=0$). The interventions on the treatment process are defined in terms of longitudinal modified treatment policies (Díaz & van der Laan, 2012; Haneuse & Rotnitzky, 2013), which are hypothetical interventions where the treatment is assigned as a new random variable $A_{t}^{\mathbbm{d}}$ (which may depend on the natural value of treatment as explained below) instead of being assigned according to the structural equation model. An intervention that sets the treatments up to time $t-1$ to $\bar{A}_{t-1}^{\mathbbm{d}}$ generates a counterfactual variable $A_{t}(\bar{A}_{t-1}^{\mathbbm{d}})$, which is referred to as the natural value of treatment (Richardson & Robins, 2013; Young et al., 2014), and represents the value of treatment that would have been observed at time $t$ under an intervention carried out up until time $t-1$ but discontinued thereafter. An intervention on all the treatment and censoring variables up to $t=\tau$ generates a counterfactual outcome $Y_{\tau+1}(\bar{A}^{\mathbbm{d}})$. Causal effects are defined as contrasts between the counterfactual probability $\mathsf{P}[Y_{\tau+1}(\bar{A}^{\mathbbm{d}})=1]$ of not experiencing a failure event by the end of the study under interventions implied by different interventions $\mathbbm{d}$, as defined below.
The causal effects we define are characterized by a user-given function $\mathbbm{d}(a_{t},h_{t},\varepsilon_{t})$ that maps a given treatment value $a_{t}$ (i.e., the “natural value”) and a history $h_{t}$ into a new treatment value. The function $\mathbbm{d}$ is also allowed to depend on a randomizer $\varepsilon_{t}$. This definition generalizes static interventions (if $\mathbbm{d}(a_{t},h_{t},\varepsilon_{t})$ is a constant), dynamic interventions (if $\mathbbm{d}(a_{t},h_{t},\varepsilon_{t})$ depends on $h_{t}$ but not on $a_{t}$ nor $\varepsilon_{t}$), and stochastic interventions (if $\mathbbm{d}(a_{t},h_{t},\varepsilon_{t})$ depends on $h_{t}$ and $\varepsilon_{t}$ but not on $a_{t}$). In what follows we will assume that the randomizer is drawn independently across units and independently of $U$, and whose distribution does not depend on $\mathsf{P}$. For fixed values $\bar{a}_{t}$, and $\bar{l}_{t}$, we recursively define $a_{t}^{\mathbbm{d}}=\mathbbm{d}(a_{t},h^{\mathbbm{d}}_{t},\varepsilon_{t})$, where $h^{\mathbbm{d}}_{t}=(\bar{a}_{t-1}^{\mathbbm{d}},\bar{l}_{t})$. The intervention is undefined if $h^{\mathbbm{d}}_{t}$ is such that $d_{s}^{\mathbbm{d}}=1$ or $y_{s}^{\mathbbm{d}}=1$ for any $s\leq t$. The LMTPs that we study are thus defined as $A^{\mathbbm{d}}_{t}=\mathbbm{d}(A_{t}(\bar{A}_{t-1}^{\mathbbm{d}}),H_{t}(\bar{A}_{t-1}^{\mathbbm{d}}),\varepsilon_{t})$ for a user-given function $\mathbbm{d}$. In what follows, we let $\mathsf{g}_{A,t}(a_{t}\mid h_{t})$ denote the density or probability mass function of $A_{t}$ conditional on $C_{t-1}=R_{t}=1$ and $H_{t}=h_{t}$, and we let $\mathsf{g}_{C,t}(a_{t},h_{t})$ denote $\mathsf{P}(C_{t}=1\mid C_{t-1}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t})$. Furthermore, we use $\mathsf{g}_{t}^{\mathbbm{d}}(a_{t}\mid h_{t})$ to denote the density or probability mass function of $\mathbbm{d}(A_{t},H_{t},\varepsilon_{t})$ conditional on $C_{t-1}=R_{t}=1$ and $H_{t}=h_{t}$.
To ground the ideas and illustrate the usefulness of this framework, we now consider several examples of regimes $\mathbbm{d}$ that yield interesting and scientifically informative causal contrasts. More examples appear in Robins et al. (2004); Young et al. (2014); Díaz et al. (2021).
Example 1 (Additive shift LMTP).
Let $A_{t}$ denote a vector of numerical treatments, such as a drug dose or energy expenditure through physical activity. Assume that $A_{t}$ is supported as $\mathsf{P}(A_{t}\leq u_{t}(h_{t})\mid H_{t}=h_{t})=1$ for some $u_{t}$. Then, for a user-given value $\delta$, we let
$\mathbbm{d}(a_{t},h_{t})=\begin{cases}a_{t}+\delta&\text{if }a_{t}+\delta\leq u_{t}(h_{t})\\ a_{t}&\text{if }a_{t}+\delta>u_{t}(h_{t}).\end{cases}$ (1)
We define $A_{t}^{\mathbbm{d}}=\mathbbm{d}(A_{t}(\bar{A}_{t-1}^{\mathbbm{d}}),H_{t}(\bar{A}_{t-1}^{\mathbbm{d}}))$. This intervention has been discussed by Díaz & van der Laan (2012), Díaz & van der Laan (2018), and Haneuse & Rotnitzky (2013) in the context of single time-point treatments; by Díaz & Hejazi (2020) and Hejazi et al. (2022c) in the context of causal mediation analysis; and by Hejazi et al. (2020) in the context of studies with two-phase sampling designs. This intervention considers hypothetical worlds in which the natural treatment at time $t$ is increased by a user-given value $\delta$, whenever such an increase is feasible for a unit with history $H_{t}(A_{t-1}^{\mathbbm{d}})$. Specifically, note that if the treatment is supported in the real numbers, the MTP could simply be defined as $\mathbbm{d}(a_{t},h_{t})=a_{t}+\delta$. However, when the treatment is supported as $\mathsf{P}(A_{t}\leq u_{t}(h_{t})\mid H_{t}=h_{t})=1$, the intervention $\mathbbm{d}(a_{t},h_{t})=a_{t}+\delta$ will not be feasible for units for which $a_{t}+\delta>u_{t}(h_{t})$. We therefore set $\mathbbm{d}(a_{t},h_{t})=a_{t}$ for those units. Note that setting $\mathbbm{d}(a_{t},h_{t})=u_{t}(h_{t})$ could also be an interesting option for these units. However, this option would lead to a parameter that violates Condition 1 below, which is necessary to achieve $\sqrt{n}-$consistent estimation.
Example 2 (Multiplicative shift LMTP).
Let $A_{t}$ denote a vector of numerical treatments such as pollutant concentrations for various pollutants. To define this intervention, assume that $A_{t}$ is supported as $\mathsf{P}(A_{t}\geq l_{t}(h_{t})\mid H_{t}=h_{t})=1$ for some $l_{t}$. Then, for a user-given value $10<\delta<$ 1, we let
$\mathbbm{d}(a_{t},h_{t})=\begin{cases}a_{t}\times\delta&\text{if }a_{t}\times\delta\geq l_{t}(h_{t})\\ a_{t}&\text{if }a_{t}\times\delta<l_{t}(h_{t}).\end{cases}$ (2)
The intervention implied by this function considers hypothetical worlds where the natural treatment at time $t$ is increased by a user-given factor $\delta$, whenever such increase is feasible for a unit with history $H_{t}(A_{t-1}^{\mathbbm{d}})$. This intervention seems useful, for example, to solve current problems in environmental epidemiology related to the effect of environmental mixtures (Gibson et al., 2019).
Example 3 (Incremental propensity score interventions based on the odds and risk ratio).
Kennedy (2019) proposed an intervention tailored to binary treatments in which the post-intervention treatment is a draw from a user-given distribution $\mathsf{g}_{t}^{\mathbbm{d}}(a_{t}\mid h_{t})$, where this distribution is chosen such that the odds ratio comparing the likelihood of treatment under intervention versus the actual treatment is a user-given value $\delta$. Specifically, consider $\varepsilon_{t}\sim U(0,1)$, and define $\mathbbm{d}(h_{t},\varepsilon_{t})=\mathds{1}\{\varepsilon_{t}<\mathsf{g}_{t}^{\mathbbm{d}}(1\mid h_{t})\}$ where we define $\mathsf{g}_{t}^{\mathbbm{d}}(1\mid h_{t})=\delta\mathsf{g}_{t}(1\mid h_{t})/\{\delta\mathsf{g}_{t}(1\mid h_{t})+\mathsf{g}_{t}(0\mid h_{t})\}$. Identifying the effect of this intervention does not require the positivity assumption. Thus, incremental propensity score interventions are very useful in settings where violations of the positivity assumption preclude identification of static effects such as the average treatment effect.
Recently, Wen et al. (2021) proposed a similar intervention that instead uses the risk ratio to quantify the likelihood of treatment under the intervention. Their intervention is defined in terms of a random draw from a distribution $\mathsf{g}_{t}^{\mathbbm{d}}(a_{t}\mid h_{t})=a_{t}\delta\mathsf{g}_{t}(1\mid h_{t})+(1-a_{t})(1-\delta\mathsf{g}_{t}(1\mid h_{t}))$ (compared to Wen et al. (2021) we have flipped the values 0 and 1 for ease of exposition). The following LMTP yields the same identifying functional as that of Wen et al. (2021):
[…]
where $\varepsilon_{t}$ is a random draw from the uniform distribution on $(0,1)$. As before, we make use of the definition $A_{t}^{\mathbbm{d}}=\mathbbm{d}(A_{t}(\bar{A}_{t-1}^{\mathbbm{d}}),H_{t}(\bar{A}_{t-1}^{\mathbbm{d}}))$. Note that the probability of treatment under the intervention is equal to $\mathsf{g}_{\delta}(1\mid h_{t})=\delta\times\mathsf{g}(1\mid h_{t})$; thus, giving a risk-ratio interpretation to $\delta$. Importantly, a unit with zero probability of treatment conditional on their history also has zero probability of treatment under the intervention. This weakens the positivity assumptions required for identification and estimation, as we will discuss below.
Example 4 (Dynamic treatment initiation strategies with grace period).
Let $A_{t}$ denote a binary treatment variable. Consider an intervention of the form “for a user-given grace period $m$, if a condition for treatment initiation is met at or before $t-m$ then start treatment at $t$; otherwise, do not intervene.” Letting $L^{\prime}_{t}$ denote an indicator that the condition has been met, this intervention can be operationalized as an LMTP as follows:
[…]
This type of intervention has been considered in the context of treatment initiation for HIV (van der Laan et al., 2005; Cain et al., 2010), where the condition $L_{t}^{\prime}$ is that CD4+ cell count drops below a pre-specified value. In this type of application, realistic treatment policies may involve a delay of treatment initiation for logistical reasons. Accordingly, allowing for a grace period quantifies an effect that is arguably of more practical relevance.
Remark 1.
Wen et al. (2021) study stochastic interventions defined as a random draw from a density that can be expressed as a mixture between a known density $\mathsf{f}_{t}(a_{t}\mid h_{t})$ and the density of the observed data $\mathsf{g}_{t}(a_{t}\mid h_{t})$. The functional identifying the mean counterfactual outcome under this stochastic intervention is the same as the functional identifying the counterfactual outcome under an LMTP defined by
$\mathbbm{d}(a_{t},h_{t},\varepsilon_{t})=\mathds{1}\{\varepsilon_{t}\leq c_{t}(h_{t})\}q_{t}+\mathds{1}\{\varepsilon_{t}>c_{t}(h_{t})\}a_{t},$
where $\varepsilon_{t}$ is a random draw from a uniform distribution in $[0,1]$, $10<c_{t}(h_{t})<$ 1 is a mixture constant, and $q_{t}$ represents a draw from $\mathsf{f}_{t}(a_{t}\mid h_{t})$. Efficient estimators for this type of LMTP have been previously proposed by Díaz et al. (2021). Wen et al. (2021) discuss the stochastic interventions version of these effects, but incorrectly claim that the estimators they develop are not particular cases of the methodology of Díaz et al. (2021).
2.1 Identification of the effect of LMTPs under competing risks
The following assumptions will be sufficient to prove identification:
A1Supported treatment intervention.
If $(a_{t},h_{t})\in\mathop{\mathrm{supp}}\{A_{t},H_{t}\}$ then $(\mathbbm{d}(a_{t},h_{t},\varepsilon_{t}),h_{t})\in\mathop{\mathrm{supp}}\{A_{t},H_{t}\}$ for $t\in\{1,\ldots,\tau\}$ and all values of $\varepsilon_{t}$.
A2Positivity of loss-to-follow-up mechanism.
$\mathsf{P}\{\mathsf{g}_{C,t}(A_{t},H_{t})>0\}=1$ for $t\in\{1,\ldots,\tau\}$.
A3Strong sequential randomization of treatment mechanism.
$U_{A,t}\mbox{$\perp\!\!\!\perp$}(\underline{U}_{D,{t+1}},\underline{U}_{Y,{t+1}},\allowbreak\underline{U}_{L,{t+1}},\allowbreak\underline{U}_{A,t+1})\mid(H_{t},R_{t}=C_{t}=1)$ for all $t\in\{1,\ldots,\tau\}$.
A4Sequential randomization of loss-to-follow-up mechanism.
$U_{C,t}\mbox{$\perp\!\!\!\perp$}(\underline{U}_{D,{t+1}},\underline{U}_{Y,t+1},\underline{U}_{L,t+1},\allowbreak\underline{U}_{A,t+1})\mid(A_{t},H_{t},R_{t}=1)$ for all $t\in\{1,\ldots,\tau\}$.
Assumptions A1 and A2 are required so that the interventions considered are supported in the data. In words, Assumption A1 requires that if there is an individual with treatment level $a_{t}$ and covariate values $h_{t}$ who is event-free and uncensored at time $t$, there must also be an individual with treatment level $\mathbbm{d}(a_{t},h_{t},\varepsilon_{t})$ and covariate values $h_{t}$ who is event-free and uncensored at time $t$. Assumption A1 would be violated if the intervention $\mathbbm{d}$ is infeasible at time $t$ for a participant with covariate history $h_{t}$. As an example of this, consider our illustrative application, letting $A_{t}$ be an indicator of having been intubated at time $t$, and letting $H_{t}$ denote a high-dimensional vector containing all the information collected for a patient up until time $t$, including information on blood oxygen saturation. Consider a modified treatment policy that would delay intubation by a number of days $\delta$ for a large value $\delta$. Note that intubation is a salvage therapy that is almost always used when a patient reaches a low oxygen saturation point. Therefore, Assumption A1 may be violated for this LMTP since a patient with very low blood oxygen saturation who is intubated at time $t$ has a very low chance of not being intubated in the vicinity of time $t$.
Assumption A2 states that for every possible value of the covariates $h_{t}$ that has positive probability of occurring, there are uncensored individuals. Assumption A3 is required to identify the effect of an LMTP. It states that, among event-free uncensored individuals, $H_{t}$ contains all the common causes of treatment at time $t$ and future events, treatments, and covariates. Assumption A4 is required to identify the effect of a static intervention that prevents loss to follow-up. It states that, among event-free uncensored individuals, $H_{t}$ contains all the common causes of censoring at time $t$ as well as future events and covariates.
Proposition 1 (Identification of the effect of LMTPs with time-to-event outcomes subject to competing events).
Set $\mathsf{q}_{\tau+1}=Y_{\tau+1}$ and $R_{\tau+1}=1$. For $t=\tau,\ldots,1$, recursively define
$\mathsf{q}_{t}:(a_{t},h_{t})\mapsto\mathsf{E}\left[R_{t+1}\times\mathsf{q}_{t+1}(A_{t+1}^{\mathbbm{d}},H_{t+1})\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\right],$ (5)
and define $\theta=\mathsf{E}[\mathsf{q}_{1}(A_{1}^{\mathbbm{d}},L_{1})]$. Under Assumptions A1-A4, $\mathsf{P}[Y_{\tau+1}(\bar{A}^{\mathbbm{d}})=1]$ is identified as $\theta$.
The identification result above, proved in the supplementary materials, extends the identification result of Young et al. (2020) to LMTPs, and the identification result of Díaz et al. (2021) to the case of competing events. This result allows researchers to estimate from data the counterfactual probability $\mathsf{P}[Y_{\tau+1}(\bar{A}^{\mathbbm{d}})=1]$, interpreted as the outcome rate in a hypothetical world where the modified treatment policy $\mathbbm{d}$ was implemented.
Next, we present efficiency theory and efficient estimators for $\theta$, based on an extension of our prior work (Díaz et al., 2021). These estimators allow the use of slow-converging data-adaptive regression techniques to obtain estimators of $\theta$ that converge at parametric rates. The general approach involves finding a von-Mises-type approximation (von Mises, 1947; van der Vaart, 1998; Robins et al., 2009; Bickel et al., 1997) for the parameter $\mathsf{q}_{L,t}$, which can be intuitively understood to be a first-order expansion with second-order error (remainder) terms. As the errors in the expansion are second-order, the resulting estimator of $\theta$ will be $\sqrt{n}$-consistent and asymptotically normal as long as the second-order error terms converge to zero at rate $\sqrt{n}$. This would be satisfied, for example, if all the regression functions used for estimation converge at rate $n^{1/4}$ (see §3). The reader interested in this general theory is encouraged to consult the works of Buckley & James (1979); Rubin & van der Laan (2007); Robins (2000); Robins et al. (1994); van der Laan & Robins (2003); Bang & Robins (2005); van der Laan & Rubin (2006); van der Laan & Rose (2011, 2018); Luedtke et al. (2017); Rotnitzky et al. (2017).
3 Efficient estimation via sequentially doubly robust regression
As noted by Díaz et al. (2021), non-parametric $\sqrt{n}$-consistent estimation is not possible for arbitrary functions $\mathbbm{d}$ when the treatment is continuous. Thus, we impose the following assumption, which is a generalization of an assumption first used by Haneuse & Rotnitzky (2013). In what follows, we will assume that the function $\mathbbm{d}$ does not depend on the distribution of the observed data $\mathsf{P}$.
1Piecewise smooth invertibility.
Assume that the cardinality of $A_{t}$ is $p$. For each $h_{t}$, assume the support of $A_{t}$ conditional on $H_{t}=h_{t}$ may be partitioned into $p$-dimensional subintervals ${\cal I}_{t,j}(h_{t}):j=1,\ldots,J_{t}(h_{t})$ such that $\mathbbm{d}(a_{t},h_{t},\varepsilon_{t})$ is equal to some $\mathbbm{d}_{j}(a_{t},h_{t},\varepsilon_{t})$ in ${\cal I}_{t,j}(h_{t})$ and $\mathbbm{d}_{j}(\cdot,h_{t},\varepsilon_{t})$ has inverse function $\mathbbm{b}_{j}(\cdot,h_{t},\varepsilon_{t})$ with Jacobian $\mathbbm{b}_{j}^{\prime}(\cdot,h_{t},\varepsilon_{t})$ with respect to $a_{t}$.
In what follows, we assume that either (i) $A_{t}$ is a discrete random variable for all $t$, or (ii) $A_{t}$ is a continuous random variable and the modified treatment policy $\mathbbm{d}$ satisfies Condition 1. Next, for continuous $A_{t}$ we define
$\mathsf{g}_{A,t}^{\mathbbm{d}}(a_{t}\mid h_{t})=\sum_{j=1}^{J_{t}(h_{t})}\int\mathds{1}_{t,j}\{\mathbbm{b}_{j}(a_{t},h_{t},\varepsilon_{t}),h_{t}\}\times\mathsf{g}_{A,t}\{\mathbbm{b}_{j}(a_{t},h_{t},\varepsilon_{t})\mid h_{t}\}\times\big{|}\det\{\mathbbm{b}_{j}^{\prime}(a_{t},h_{t},\varepsilon_{t})\}\big{|}\,\mathrm{d}\mathsf{P}(\varepsilon_{t}),$ (6)
where $\mathds{1}_{t,j}\{u,h_{t}\}=1$ if $u\in{\cal I}_{t,j}(h_{t})$ and $\mathds{1}_{t,j}\{u,h_{t}\}=0$ otherwise. Under Condition 1, application of the formula for the density of a transformation shows that the p.d.f. of $A_{t}^{\mathbbm{d}}$ conditional on the history $h_{t}$ is $\mathsf{g}_{A,t}^{\mathbbm{d}}(a_{t}\mid h_{t})$. In the case of Example 2, the post-intervention p.d.f. becomes
$\mathsf{g}_{A,t}^{\mathbbm{d}}(a_{t}\mid h_{t})=\mathsf{g}_{A,t}(a_{t}\mid h_{t})\mathds{1}\{a_{t}\delta<l_{t}(h_{t})\}+\delta^{-1}\mathsf{g}_{A,t}(a_{t}\delta^{-1}\mid h_{t})\mathds{1}\{a_{t}\geq l_{t}(h_{t})\},$
which shows that piecewise smoothness is sufficient to handle interventions such as (2) which are not smooth in the full range of the treatment. Condition 1 and expression (6) ensure that we can use the change of variable formula when computing integrals over $\mathcal{A}_{t}$ for continuous treatments. This is useful for studying properties of the parameter and estimators we propose.
For discrete treatments, we let
$\mathsf{g}_{A,t}^{\mathbbm{d}}(a_{t}\mid h_{t})=\sum_{s_{t}\in\mathcal{A}_{t}}\int\mathds{1}\{\mathbbm{d}(s_{t},h_{t},\varepsilon_{t})=a_{t}\}\mathsf{g}_{A,t}(s_{t}\mid h_{t})\,\mathrm{d}\mathsf{P}(\varepsilon_{t}).$ (7)
Evaluated using the definition of the incremental propensity score intervention based on the risk ratio given in Example 3, this expression yields
$\mathsf{g}_{A,t}^{\mathbbm{d}}(a_{t}\mid h_{t})=a_{t}\delta\mathsf{g}_{A,t}(1\mid h_{t})+(1-a_{t})[1-\delta\mathsf{g}_{A,t}(1\mid h_{t})],$
which clarifies that the risk ratio for receiving the treatment under the intervention versus the observed data is equal to $\delta$. In what follows, it will be useful to define the parameter
$\mathsf{w}_{t}(c_{t},a_{t},h_{t})=\frac{\mathsf{g}_{A,t}^{\mathbbm{d}}(a_{t}\mid h_{t})}{\mathsf{g}_{A,t}(a_{t}\mid h_{t})}\frac{c_{t}\times r_{t}}{\mathsf{g}_{C,t}(a_{t},h_{t})},$
and the functions
$\varphi_{t}:x\mapsto\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{w}_{k}(c_{k},a_{k},h_{k})\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{d}},h_{s+1})-\mathsf{q}_{s}(a_{s},h_{s})\}+\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})$ (8)
for $t=\tau,\ldots,1$ which map $x=(d_{t},y_{t},l_{t},a_{t},c_{t}:t\in\{1,\ldots,\tau\})$ to the real line. When necessary, we use the notation $\varphi_{t}(x;\eta)$ or $\varphi_{t}(x;\underline{\eta}_{t})$ to highlight the dependence of $\varphi_{t}$ on $\underline{\eta}_{t}=(\mathsf{w}_{t},\mathsf{q}_{t},\ldots,\mathsf{w}_{\tau},\mathsf{q}_{\tau})$ We also use $\eta$ to denote $(\mathsf{w}_{1},\mathsf{q}_{1},\ldots,\mathsf{w}_{\tau},\mathsf{q}_{\tau})$, and define $\varphi_{\tau+1}(X;\eta)=Y$.
Estimation of $\theta$ will proceed by regression of the transformation $\varphi_{t+1}$ sequentially from $t=\tau$ to $0=1$ to obtain estimates $\mathsf{q}_{t}$. We say that $\varphi_{t+1}$ is a sequentially doubly robust unbiased transformation for $\mathsf{q}_{t}$ due to the following proposition:
Proposition 2 (Sequentially doubly robust transformation).
Let $\eta^{\prime}$ be such that either $\mathsf{q}^{\prime}_{s}=\mathsf{q}_{s}$ or $\mathsf{w}^{\prime}_{s}=\mathsf{w}_{s}$ for all $s>t$. Then, in the event $R_{t}=1$, we have
$\mathsf{E}\big{[}R_{t+1}\times\varphi_{t+1}(X;\eta^{\prime})\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]}=\mathsf{q}_{t}(a_{t},h_{t}).$
This proposition is a straightforward application of the von-Mises expansion given in Lemma 1 in the supplementary materials. Another consequence of Lemma 1 is that the efficient influence function for estimating $\theta=\mathsf{E}[\mathsf{q}_{1}(A^{\mathbbm{d}},L_{1})]$ in the non-parametric model is given by $\varphi_{1}(X)-\theta$ (see Díaz et al., 2021). Proposition 2 motivates the construction of the sequential regression estimator by iteratively regressing an estimate of the data transformation $\varphi_{t+1}(X;\eta)$ on $(A_{t},H_{t})$ among individuals with $C_{t-1}=R_{t}=1$, starting at $\varphi_{\tau+1}(X;\eta)=Y$.
Remark 2 (Double robustness of incremental propensity score interventions).
Consider the odds ratio incremental propensity score interventions of Example 3, where we let $\varepsilon_{t}\sim U(0,1)$, and define $\mathbbm{d}(h_{t},\varepsilon_{t})=\mathds{1}\{\varepsilon_{t}<\mathsf{g}_{t}^{\mathbbm{d}}(1\mid h_{t})\}$. The result in Proposition 2 also holds for this intervention, but it does not allow for doubly robust estimation. To see why, notice that $\mathsf{q}_{t}$ itelse depends on $\{\mathsf{g}_{s}:s>t\}$ (or equivalently, on $\{\mathsf{w}_{s}:s>t\}$) through $\mathbbm{d}$. Therefore, if $\mathsf{q}^{\prime}_{s}=\mathsf{q}_{s}$ and $\mathsf{w}^{\prime}_{s}\neq\mathsf{w}_{s}$ for all $s>t$, we get $\mathsf{E}\big{[}R_{t+1}\times\varphi_{t+1}(X;\eta^{\prime})\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]}=\mathsf{q}_{t}(a_{t},h_{t})$, where $\mathsf{q}_{t}$ is evaluated at the misspecified $\{\mathsf{w}_{s}^{\prime}:s>t\}$. This is aligned with the findings of Kennedy (2019). In contrast, the risk ratio incremental propensity score intervention does allow for doubly robust estimation, since $\mathbbm{d}_{t}$ does not depend on $\mathsf{w}_{t}$.
In order to avoid imposing entropy conditions on the initial estimators, we use sample splitting and cross-fitting (Klaassen, 1987; Zheng & van der Laan, 2011; Chernozhukov et al., 2018). Let ${\cal V}_{1},\ldots,{\cal V}_{J}$ denote a random partition of the index set $\{1,\ldots,n\}$ into $J$ prediction sets of approximately the same size. That is, ${\cal V}_{j}\subset\{1,\ldots,n\}$; $\bigcup_{j=1}^{J}{\cal V}_{j}=\{1,\ldots,n\}$; and ${\cal V}_{j}\cap{\cal V}_{j^{\prime}}=\emptyset$. In addition, for each $j$, the associated training sample is given by ${\cal T}_{j}=\{1,\ldots,n\}\setminus{\cal V}_{j}$. We let $\hat{\eta}_{j}$ denote the estimator of $\eta$ obtained by training the corresponding prediction algorithm using only data in the sample ${\cal T}_{j}$. Further, we let $j(i)$ denote the index of the validation set containing unit $i$. For preliminary estimates $\hat{\mathsf{w}}_{1,j(i)},\ldots,\hat{\mathsf{w}}_{\tau,j(i)}$, the estimator is defined as follows:
  1. Step 1:
    Initialize $\varphi_{\tau+1}(X_{i};\underline{\check{\eta}}_{\tau,j(i)})=Y_{i}$ for $i=1,\ldots,n$.
  2. Step 2:
    For $t=\tau,\ldots,1$:
    • Compute the pseudo-outcome $\check{Y}_{t+1,i}=R_{t+1}\times\varphi_{t+1}(X_{i};\underline{\check{\eta}}_{t,j(i)})$ for all $i=1,\ldots,n$.
    • For $j=1,\ldots,J$:
      • Regress $\check{Y}_{t+1,i}$ on $(A_{t,i,}H_{t,i})$ using any regression technique and using only data points $i\in\mathcal{T}_{j}$.
      • Let $\check{\mathsf{q}}_{t,j}$ denote the output of the above regression, update $\underline{\check{\eta}}_{t,j}=(\hat{\mathsf{w}}_{t,j},\check{\mathsf{q}}_{t,j},\ldots,\hat{\mathsf{w}}_{\tau,j},\check{\mathsf{q}}_{\tau,j})$, and iterate.
  3. Step 3:
    Define the sequential regression estimator as
    $\hat{\theta}=\frac{1}{n}\sum_{i=1}^{n}\varphi_{1}(X_{i};\check{\eta}_{j(i)}).$
Implementation of the above estimator requires an algorithm for estimation of the parameters $\mathsf{w}_{t}$ for all $t$. This parameter may be estimated component-wise as follows. First, the censoring mechanism $\mathsf{g}_{C,t}$ may be estimated through any regression or classification procedure by regressing the indicator $C_{t}$ on data $(A_{t},H_{t})$ among individuals with $R_{t}=C_{t}=1$. Thus, it only remains to estimate the density ratio $\mathsf{g}_{t}^{\mathbbm{d}}(a_{t}\mid h_{t})/\mathsf{g}_{t}(a_{t}\mid h_{t})$. An option to estimate this density ratio is to first obtain an estimate of the density $\mathsf{g}_{t}(a_{t}\mid h_{t})$, and then to obtain an estimator of $\mathsf{g}_{t}^{\mathbbm{d}}(a_{t}\mid h_{t})$ by applying formula (6) or (7). This approach is problematic in the case of continuous or multivariate treatments due both to the curse of dimensionality and to the fact that flexible and computationally efficient estimation of conditional densities is an under-developed area of machine learning (relative to regression), though some suitable techniques do exist (e.g., Díaz & van der Laan, 2011; Hejazi et al., 2022b, a). This approach can also be problematic for discrete treatments of high cardinality. Since the conditional densities themselves are unnecessary for estimation, and direct estimation of the ratio $\mathsf{g}_{t}^{\mathbbm{d}}(a_{t}\mid h_{t})/\mathsf{g}_{t}(a_{t}\mid h_{t})$ suffices, we propose to use Díaz et al. (2021)’s approach to direct estimation of density ratios. Briefly, this approach starts by constructing an augmented artificial dataset of size $2n$ in which each observation is duplicated. For each observation, one duplicate gets assigned the post-intervention treatment value $\mathbbm{d}(A_{t},H_{t})$ and the other gets the actual observed value $A_{t}$. Then, the task of density ratio estimation reduces to the task of classifying which record belongs to the intervened-upon unit, based on the treatment and the history $H_{t}$. Specifically, Díaz et al. (2021) show that the odds of this classification problem is equal to the density ratio. Importantly, data-splitting procedures such as cross-validation or cross-fitting must be performed such that the two records from the same participants are always kept in the same fold.
The above estimation algorithm requires the implementation of several regression procedures. Many algorithms from the machine learning literature are available for such regression problems, including those based on ensembles of regression trees, splines, etc. As stated in the asymptotic normality result below, it is important that the regression procedures used approximate as well as possible the relevant components of the true data-generating mechanism. Increasing the predictive performance of these regression fits can entail estimator selection from a large list of candidate estimators. If the sample size is large enough, a principled solution to this problem is the use of Super Learning (van der Laan et al., 2006, 2007), a form of model stacking (Wolpert, 1992; Breiman, 1996) that constructs an ensemble model as a convex combination of candidate algorithms from a user-supplied regression library, where the weights in the combination are computed based on cross-validation and are such that they minimize the holdout prediction error of the resulting estimator.
The asymptotic distribution of $\theta$ is Gaussian with variance equal to the efficiency bound, under conditions. To state the conditions, we first define the data-dependent parameter
$\check{\mathsf{q}}_{t}^{\dagger}(a_{t},h_{t})=\mathsf{E}\big{[}R_{t+1}\varphi_{t+1}(X;\underline{\check{\eta}}_{t})\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]},$
where the outer expectation is only with respect to the distribution $\mathsf{P}$ of $X$ (i.e., $\check{\eta}$ is fixed). The following theorem is a consequence of Theorem 4 of Díaz et al. (2021):
Theorem 1 (Consistency and weak convergence of SDR estimator).
We have
  1. (i)
    If, for each time $t$, either $||\hat{\mathsf{w}}_{t}-\mathsf{w}_{t}||=o_{\mathsf{P}}(1)$ or $||\check{\mathsf{q}}_{t}-\check{\mathsf{q}}^{\dagger}_{t}||=o_{\mathsf{P}}(1)$, then $\hat{\theta}=\theta+o_{\mathsf{P}}(1)$.
  2. (ii)
    If $\sum_{t=1}^{\tau}||\hat{\mathsf{w}}_{t}-\mathsf{w}_{t}||\,||\check{\mathsf{q}}_{t}-\check{\mathsf{q}}_{t}^{\dagger}||=o_{\mathsf{P}}(n^{-1/2})$ and $\mathsf{P}\{\mathsf{w}_{t}(A_{t},H_{t})<c\}=\mathsf{P}\{\hat{\mathsf{w}}_{t}(A_{t},H_{t})<c\}=1$ for some $c<\infty$, then
    $\sqrt{n}(\hat{\theta}-\theta)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varphi_{1}(X_{i};\eta)+o_{P}(1),$
    implying that $\sqrt{n}(\hat{\theta}-\theta)\rightsquigarrow N(0,\sigma^{2})$, where $\sigma^{2}=\operatorname{\mathsf{Var}}\{\varphi_{1}(X;\eta)\}$ is the non-parametric efficiency bound.
Note that this means that $\hat{\theta}$ is sequentially doubly robust in the sense that it is consistent if at least one of the two nuisance parameters is consistently estimated at each time point. This property is also known in the literature as $2^{\tau}$-multiple robustness. We prefer the term sequentially doubly robust for two reasons. First, contrasted to the double robustness property which requires one out of two nuisance estimators to be consistent, the term $2^{\tau}$-multiple robustness erroneously conveys the message that consistency would follow from one out of $2^{\tau}$ nuisance estimators being consistent. Second, retaining the term “double” in this description conveys a fundamental property of the estimation procedure: it is based on an expansion (given in Lemma 1 in the supplementary material) with second-order remainder terms. In this nomenclature, a triply robust estimating procedure would be one based on an expansion with third-order remainder terms, and so forth. [^1] Some of these points were brought to our attention by Edward H. Kennedy in a conversation on Twitter (see this thread, https://twitter.com/edwardhkennedy/status/1446952129218949128?s=20&t=8o_-20szw0PPfqocYiw7Dg).
A drawback of the above estimation approach is that there is no guarantee that the parameter will remain within the bounds of the parameter space. That is, it is possible, in principle, that the sequential regression estimator will provide time-to-event probabilities outside of the closed unit interval $[0,1]$. In the supplementary materials, we present a targeted minimum loss estimator (TMLE) capable of producing estimates that always lie within the unit interval. However, boundedness of the TMLE comes with a robustness tradeoff. In particular, the robustness properties of the TMLE are inferior to those of the sequentially doubly robust estimator. Specifically, consistency of the TMLE requires that either $\lVert\hat{\mathsf{w}}_{t}-\mathsf{w}_{t}\rVert=o_{\mathsf{P}}(1)$ or $\lVert\hat{\mathsf{q}}_{t}-\mathsf{q}_{t}\rVert=o_{\mathsf{P}}(1)$, for an estimator $\hat{\mathsf{q}}_{t}$ constructed based on sequential regression using (5). As a result, consistent estimation of $\mathsf{q}_{t}$ in the TMLE procedure requires consistent estimation of $\mathsf{q}_{s}:s>t$. By contrast, the sequentially doubly robust estimator provides consistent estimation of $\mathsf{q}_{t}$ when either $\mathsf{q}_{s}$ or $\mathsf{w}_{s}$ is consistently estimated for all $s>t$. Both Luedtke et al. (2017) and Díaz et al. (2021) provide further in-depth discussion on this topic. Furthermore, the conditions required for $\sqrt{n}$-consistency for this TMLE algorithm may be stronger than those required for the SDR estimator. In the supplementary materials, we provide an argument that the second-order term associated to the TMLE is bounded above by the term
$\sum_{t=1}^{\tau}\left(\lVert\hat{\mathsf{w}}_{t}-\mathsf{w}_{t}\rVert\sum_{s=t}^{\tau}\lVert\hat{\mathsf{q}}_{s}-\mathsf{q}_{s}^{\star}\rVert\right),$
where $\mathsf{q}_{t}^{\star}$ is the data-dependent parameter defined as
$\mathsf{q}_{t}^{\star}(a_{t},h_{t})=\mathsf{E}\big{[}R_{t+1}\hat{\mathsf{q}}_{t+1}(a_{t+1},h_{t+1})\mid C_{t}=R_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\big{]}.$
In comparison to the assumption in statement (ii) of the above theorem, convergence of the second-order term associated to the TMLE seems to require stronger assumptions that would also require convergence of all the cross-product terms $\lVert\hat{\mathsf{w}}_{t}-\mathsf{w}_{t}\rVert\lVert\hat{\mathsf{q}}_{s}-\mathsf{q}_{s}^{\star}\rVert$ for $s\geq t$.
Neither the SDR nor the TMLE guarantee that a survival function estimated at several points in time will result in monotonic decreasing estimates. In order to leverage the desirable robustness properties of the SDR estimator while endowing it with the properties of boundedness and monotonicity, we leverage a set of techniques proposed by Westling et al. (2020). Under their approach, out-of-bounds estimates are truncated to remain within bounds, and the possibly non-monotonic estimate of the survival curve, alongside any confidence region limits, is projected onto the space of monotonic functions by way of isotonic regression. Importantly, Westling et al. (2020) show that the resulting projected estimator retains the same asymptotic properties as the original estimator. In our case, this implies that the projected estimator retains the properties of asymptotic normality and consistency under the assumptions of Theorem 1. Specifically, simultaneous confidence bands centered around the isotonic projection constructed using the multivariate distribution implied by Theorem 1 will have correct asymptotic coverage under the assumptions of the theorem.
Under the conditions of the above theorem, a Wald-type procedure, in which the standard error may be estimated based on the empirical variance of $\varphi_{1}(X_{i};\check{\eta}_{j(i)})$, yields asymptotically valid confidence intervals and hypothesis tests. Furthermore, the asymptotic linearity of Theorem 1 implies that estimates of survival functions for multiple time points asymptotically converge to a multivariate normal distribution, which can be used to construct simultaneous confidence intervals using the method described, e.g., by Hothorn et al. (2008).
4 Illustrative application
4.1 Background on motivating study and data
To illustrate the proposed methods, we aim to answer the following question: what is the effect of invasive mechanical ventilation (IMV) on acute kidney injury (AKI) among COVID-19 patients? This question has been recently identified by an expert panel on lung–kidney interactions in critically ill patients (Joannidis et al., 2020) as an area in need of further research. AKI is a common condition in the ICU, complicating about 30% of ICU admissions, and causing increased risk of in-hospital mortality and long-term morbidity and mortality (Kes & Jukić, 2010). AKI is often viewed as a tolerable consequence of interventions to support other failing organs, such as IMV (Husain-Syed et al., 2016).
To answer this question, we will use data on approximately 3,300 patients hospitalized with COVID-19 at the New York Presbyterian Cornell, Queens, or Lower Manhattan hospitals. Our analysis focuses on patients hospitalized between March 3rd and May 15th, 2020, who did not have a previous history of Chronic Kidney Disease (CKD). This choice of timeframe is rooted in there being a lack of treatment guidelines in the initial months of the COVID-19 pandemic, which resulted in a large variability in provider practice regarding mechanical ventilation. This clinical heterogeneity in the deployment of IMV makes the problem particularly well-suited to assessment via the causal effects of LMTPs. The analytical dataset was built from two distinct databases. Demographics, comorbidity, intubation, death, and discharge data were abstracted from electronic health records by trained medical professionals into a secure RedCAP database (Goyal et al., 2020). These data were supplemented with the Weill Cornell Critical carE Database for Advanced Research (WC-CEDAR), a comprehensive data repository housing laboratory, procedure, diagnosis, medication, microbiology, and flowsheet data documented as part of standard care (Schenck et al., 2021). This dataset is unique in that it contains a large sample of patients treated in the same hospital system within a short period of time, while still exhibiting extensive treatment variability.
The most frequent reason for morbidity and mortality among patients with COVID-19 is pneumonia leading to acute respiratory distress syndrome (ARDS). Respiratory support via devices such as nasal cannulae, face masks, or endotracheal tubes (i.e., intubation and subsequent mechanical ventilation) is a primary treatment for ARDS (Hasan et al., 2020). For severe ARDS patients, an important decision must be made regarding the optimal timing of intubation based on clinical factors such as oxygen saturation, dyspnea, respiratory rate, chest radiograph, and, occasionally, external resource considerations such as ventilator or provider availability (Tobin, 2020; Thomson & Calligaro, 2021). At the outset of the COVID-19 pandemic, multiple international guidelines recommended early intubation in an effort to protect healthcare workers from infection and to avoid complications stemming from “crash” intubations (Papoutsi et al., 2021). Yet, as the pandemic progressed and multiple reports documented high mortality for mechanically ventilated patients, guidance changed to delaying intubation with the rationale that IMV increases risk of secondary infections, ventilator-induced lung injury, damage to other organs, and death (Bavishi et al., 2021; Tobin, 2006). The lungs and kidneys are physiologically closely tied, and it has been hypothesized that mechanical ventilation may cause AKI in COVID-19 patients via oxygen toxicity and capillary endothelial damage leading to inflammation, hypotension, and sepsis (Durdevic et al., 2020).
Despite these hypotheses, a causal link between intubation and AKI in patients with COVID-19 has not been definitively established. Furthermore, such a link is difficult to study in a randomized fashion due to the clinically subjective nature of intubation assignment — that is, it is impossible to pick an oxygen saturation breakpoint to safely intubate across all patients and hospital settings (Tobin et al., 2020; Perkins et al., 2020).
While the original dataset cannot be made publicly available due to patient privacy considerations, we have created a synthetic dataset that mimics the general structure of the real clinical dataset, and we have made this synthetic dataset available publicly. The R code and the synthetic dataset are available in the GitHub repository at https://github.com/kathoffman/lida-comprisks.
4.2 Time-varying variables and modified treatment policy
For each patient, study time begins on the day of hospitalization. The treatment of interest is daily maximum respiratory support, which can take three levels: no supplemental oxygen, supplemental oxygen not including IMV, and IMV. Our goal is to estimate the overall causal effect on daily AKI rates of an intervention that would delay intubation for IMV by a single day among 449 patients who received IMV for more than one day, and a prevention of IMV for patients who received IMV for exactly one day. To simplify the presentation, in what follows we refer to this intervention as a “one-day delay in intubation.” Death by other causes is treated as a competing risk for AKI development. In notation, this LMTP may be expressed as follows: Let $A_{t}\in\{0,1,2\}$, where 00 indicates no supplemental oxygen at end of day $t$, $1$ indicates supplemental oxygen not including IMV, and $2$ indicates IMV. Consider the following intervention:
$\mathbbm{d}_{t}(a_{t},h_{t})=\begin{cases}1&\text{ if }a_{t}=2\text{ and }a_{s}\leq 1\text{ for all }s<t,\\ a_{t}&\text{ otherwise.}\end{cases}$ (9)
Under this intervention, a patient who is first intubated at day $t$ would receive non-invasive oxygen support at day $t$. Otherwise, the patient would receive the oxygen support actually (i.e., “naturally”) observed at day $t$. This intervention assesses what would have happened in a hypothetical world where intubation procedures were more conservative (operationalized as a one-day delay) than they were at the beginning of the pandemic in New York City’s surge conditions, when limited information and treatments were available for COVID-19. Figure 1 shows the observed supplemental oxygen levels and observed AKI and death rates across days 1-14.
Baseline confounders include age, sex, race, ethnicity, body mass index (BMI), hospital location, home oxygen status, and comorbidities (e.g., hypertension, history of stroke, Diabetes Mellitus, Coronary Artery Disease, active cancer, cirrhosis, asthma, Chronic Obstructive Pulmonary Disease, Interstitial Lung Disease, HIV infection, and immunosuppression), and are included in $L_{1}$. Time-dependent confounders include vital signs (e.g., highest and lowest respiratory rate, oxygen saturation, temperature, heart rate, and blood pressure), laboratory results (e.g., C-Reactive Protein, BUN Creatinine Ratio, Activated Partial Thromboplastin time, Creatinine, lymphocytes, neutrophils, bilirubin, platelets, D-dimer, glucose, arterial partial pressure of oxygen, and arterial partial pressure of carbon dioxide), and an indicator for daily corticosteroid administration greater than 0.5 mg/kg body weight. Time-dependent confounders were measured irregularly for each patient. We bin them in 24-hour intervals from the time of hospitalization. In cases of multiple measurements in a 24-hour window, the clinically worst measure was used. In cases of missing baseline confounders, multiple imputation using chained equations (MICE) (van Buuren & Groothuis-Oudshoorn, 2011) is used. For missing data at later time points, the last observation is carried forward. Patients are censored at their day of hospital discharge, as AKI and vital stataus were unknown after this point. A distribution of the censoring weights is given in Figure 3 in the supplementary materials.
A possible limitation of our setup is that we treat discharge as a censoring event. If patients who remain in-hospital at the end of follow-up are very different from those who are discharged, and if such differences are not explained by measured covariates, then our results may be biased. Specifically, our analyses rely on Assumption A4, which applied to this problem states that $(A_{t},H_{t})$ contains all the common causes of discharge at time $t$ and future variables measured at times $s>t$, including future outcomes. This assumption could be violated if there is an unmeasured common cause of discharge and outcomes such as an underlying biological or clinical process not captured by the covariates listed in the above paragraph.
Figure 1: A flow diagram, or “alluvial plot”, showing the movement of patients between exposure and outcome statuses. The $x$ axis shows study day and the $y$ axis shows counts of patients. Colors indicate exposure status for patients still in the study, or outcome status for patients who have already experienced the outcome, competing risk, or loss-to-follow-up. Darker-colored bars indicate the number of patients in each exposure or outcome at that time point, while lighter-colored bars visualize patients moving between time points. The exposure contains three levels: no supplemental oxygen, supplemental oxygen not including IMV, and IMV. The outcome of interest is AKI, and death is a competing risk. Patients who are discharged are right-censored. There is a one-day lag (i.e. no exposure status shown for the last study day) so that all data used in the analysis can be visualized on a single plot.
4.3 Results
We estimate the effect of the LMTP with the SDR estimator (on account of its enhanced robustness relative to TMLE), using a version of the lmtp R package (Williams et al., 2022) modified to support the competing risks setting. The nuisance functions $\mathsf{w}_{t}$ and $\mathsf{q}_{t}$ are estimated using the Super Learner ensemble modeling algorithm (van der Laan et al., 2007) via the sl3 R package (Coyle et al., 2022). The candidate library considered by the Super Learner included a total of 32 learning algorithms, inclusive of hyperparameter variations. The algorithms included multivariate adaptive regression splines (MARS; Friedman et al., 1991); random forests (Breiman, 2001; Wright & Ziegler, 2017); extreme gradient boosted trees (Friedman, 2001; Chen & Guestrin, 2016); lasso, ridge, and elastic net (with equal weights, i.e., $\alpha=0.5$) regularized regression (Friedman et al., 2022); main terms generalized linear models (McCullagh & Nelder, 1989; Enea, 2009); generalized linear models with Bayesian priors on parameter estimates (Gelman & Hill, 2006); and an unadjusted outcome mean. The estimators of $\mathsf{w}_{t}$ and $\mathsf{q}_{t}$ were computed under a Markov assumption with lag of two days to reflect the collection of laboratory results at minimum 48-hour intervals.
Note that a complete clinical understanding of the effect of IMV would require estimating its effect on death, as these interventions are meant to save lives. Figure 2 displays the SDR-based incidence difference between the two treatment conditions for both outcomes, together with 95% simultaneous confidence bands.
Figure 2: Difference between the two treatment policies for two outcomes: AKI and mortality. 95% simultaneous confidence bands cover the sets of point estimates.
Inspection of Figure 2 makes apparent the protective effect of the delay-in-intubation policy against AKI. The estimates of the incidence difference are consistently negative and decrease steadily in time; moreover, the simultaneous confidence bands do not overlap with the null value of zero except for on the first day. Hypothesis testing of the incidence difference yields adjusted p-values $<0.001$ for all days after the first, indicating a statistically significant protective effect of the LMTP; the Bonferroni correction was used to adjust for testing multiplicity. Based on these estimates, the delay-in-intubation policy would be expected to result in a decrease in 14-day AKI rates by approximately 3.5% (simultaneous 95% CI: [2.7%, 4.3%]) relative to no intervention. Importantly, while IMV does indeed have the effect of increasing survival, its negative effect on AKI is much larger than its positive effect on survival. This raises the question of whether there are patients for whom there is no survival benefit who are being harmed by the intervention. Developing methods for identifying those patients is the subject of future methodological research.
5 Discussion
Modified treatment policies represent a recent but powerful advance in modern causal inference. These intervention regimes have a variety of applications, including (i) flexibility in the definition of effects whose identification relies on a positivity assumption that may be satisfied by design; (ii) flexibility in the definition and estimation of the effects of continuous, time-valued, and multivariate treatments; and (iii) definition of effects that depend on the natural value of treatment. Here, we have advanced the LMTP methodology, extending it to the case of survival outcomes with competing events. We have presented identification assumptions for the causal parameters, as well as a sequentially doubly robust and efficient estimation algorithm capable of leveraging the many data-adaptive regression procedures available for nuisance parameter estimation. We ensure that the estimates of the survival function remain within the parameter space (monotonic decreasing and bounded in the closed unit interval) by leveraging the monotonic projection framework of Westling & Carone (2020) and Westling et al. (2020). We have illustrated the scientific utility of our approach through application to the study of negative side effects of invasive mechanical ventilation on acute kidney injury in patients hospitalized with COVID-19, for which death by other causes serves as a competing event.
Our Proposition 2 shows that doubly robust estimation is possible for any intervention that is defined in terms of an LMTP satisfying the stated conditions. Namely, any randomized intervention defined by a known function $\mathbbm{d}(a_{t},h_{t},\varepsilon_{t})$ through either an LMTP or a general stochastic intervention accommodates doubly robust estimation, provided that: (1) either (i) $A_{t}$ is a discrete random variable for all $t$, or (ii) $A_{t}$ is a continuous random variable and the modified treatment policy $\mathbbm{d}$ satisfies Condition 1, and (2) the randomizer $\varepsilon_{t}$ is drawn independently across units and independently of $U$, and its distribution does not depend on $\mathsf{P}$. Notably, this rich class contains all interventions considered by Theorem 2 of Wen et al. (2021).
A limitation of the approach taken in this paper is that we considered data collected in discrete time. While the assumption of discrete time is sensible in many applications (e.g., most clinical measures are recorded in discrete time, whether the scale is seconds, minutes, hours, or days), our methods may be of limited applicability when the time scale is truly continuous or very fine (e.g., seconds) such that many time points need to be considered. In such cases, it may be possible to coarsen the time scale into discrete components. Whether such an approach provides sensible answers and the appropriate level of coarsening needs to be assessed on a case-by-case basis based on subject-matter relevance for the application at hand.
Conflict of Interest
The authors declare that they have no conflict of interest.
Appendix A Auxiliary results
The proofs of results in the paper rely heavily on existing identification and optimality results for general longitudinal modified treatment policies (Díaz et al., 2021). We first present those results. Consider a data structure $Z=(W_{1},E_{1},\ldots,W_{\tau},E_{\tau},Y)$. We assume these data are generated by a non-parametric structural equation model with exogenous errors $U_{W_{t}}$ and $U_{E_{t}}$ as detailed by Díaz et al. (2021). Let $\mathsf{m}_{C,\tau+1}=Y$, $V_{t}=(\bar{W}_{t},\bar{E}_{t-1})$, and let $\mathbbm{h}(e_{t},v_{t})$ denote a modified treatment policy on $E$, with $E_{t}^{\mathbbm{h}}=\mathbbm{h}(E_{t}(\bar{E}_{t-1}^{\mathbbm{h}}),V_{t}(\bar{E}_{t-1}^{\mathbbm{h}}))$ denoting the intervened exposure variable. For $t=\tau,\ldots,0$, let
$\mathsf{m}_{t}(e_{t},v_{t})=\mathsf{E}[\mathsf{m}_{t+1}(E_{t+1}^{\mathbbm{h}},V_{t+1})\mid E_{t}=e_{t},V_{t}=v_{t}].$
Furthermore, let $\mathsf{g}_{t}(e_{t}\mid v_{t})$ denote the probability density function of $E_{t}$ conditional on $V_{t}=v_{t}$, $\mathsf{g}_{t}^{\mathbbm{h}}(e_{t}\mid v_{t})$ the probability density function of $E_{t}^{\mathbbm{h}}$ conditional on $V_{t}=v_{t}$, and define
$\mathsf{z}_{t}(e_{t},v_{t})=\frac{\mathsf{g}_{t}^{\mathbbm{h}}(e_{t}\mid v_{t})}{\mathsf{g}_{t}(e_{t}\mid v_{t})}.$
Define
$\psi_{t}(z)=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})\right)\{\mathsf{m}_{s+1}(e_{s+1}^{\mathbbm{h}},v_{s+1})-\mathsf{m}_{s}(e_{s},v_{s})\}+\mathsf{m}_{t}(e_{t}^{\mathbbm{h}},v_{t}).$
Then Díaz et al. (2021) proved the following three results.
Theorem 2.
Let $Y(\bar{E}^{\mathbbm{h}})$ denote a counterfactual outcome in a hypothetical world where $\bar{E}=\bar{E}^{\mathbbm{h}}$. Assume $U_{E,t}\mbox{$\perp\!\!\!\perp$}\underaccent{\bar}{U}_{W,t+1}\mid V_{t}$ and that $(e_{t},v_{t})\in\mathop{\mathrm{supp}}\{E_{t},V_{t}\}$ implies $(\mathbbm{h}(e_{t},v_{t}),v_{t})\in\mathop{\mathrm{supp}}\{E_{t},V_{t}\}$. Then $\lambda=\mathsf{E}[\mathsf{m}_{1}(E_{1}^{\mathbbm{h}},V_{1})]$ identifies $\mathsf{E}\{Y(\bar{E}^{\mathbbm{h}})\}$.
Theorem 3.
The efficient influence function for estimating $\lambda$ in the non-parametric model is given by $\psi_{1}(Z)-\lambda$.
Theorem 4.
For any fixed sequence $(\mathsf{m}_{1}^{\prime},\mathsf{z}_{1}^{\prime},\ldots,\mathsf{m}_{\tau}^{\prime},\mathsf{z}_{\tau}^{\prime})$, define
$\mathsf{K}_{t,s}^{\prime}=\prod_{l=t+1}^{s-1}\mathsf{z}_{l}^{\prime}(E_{l},V_{l}),$
and
$\mathsf{U}_{t}(e_{t},v_{t})=\sum_{s=t+1}^{\tau}\mathsf{E}\left[\mathsf{K}_{t,s}^{\prime}\{\mathsf{z}_{s}^{\prime}(E_{s},V_{s})-\mathsf{z}_{s}(E_{s},V_{s})\}\{\mathsf{m}_{s}^{\prime}(E_{s},V_{s})-\mathsf{m}_{s}(E_{s},V_{s})\}\mid E_{t}=e_{t},V_{t}=v_{t}\right].$
Then we have
$\mathsf{m}_{t}(e_{t},v_{t})=\mathsf{E}[\psi_{t+1}^{\prime}(Z)\mid E_{t}=e_{t},V_{t}=v_{t}]+\mathsf{U}_{t}(e_{t},v_{t})$
where $\psi_{t+1}^{\prime}$ is defined as above with $\mathsf{m}_{s}$ and $\mathsf{z}_{s}$ replaced by $\mathsf{m}_{s}^{\prime}$ and $\mathsf{z}_{s}^{\prime}$, respectively.
Appendix B Identification (Proposition 1)
Proof Let $V_{t}=(D_{t},Y_{t},L_{t})$ and $E_{t}=(A_{t},C_{t})$. Let $\mathbbm{h}(E_{t},V_{t})=(\mathbbm{d}(A_{t},H_{t}),1)$. This theorem follows from application of Theorem 2 above. First, note that for $t=\tau$, we have
$\mathsf{m}_{t}(e_{t}^{\mathbbm{h}},v_{t})$ $=\mathsf{E}[Y_{\tau+1}\mid D_{t}=d_{t},Y_{t}=y_{t},H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}},C_{t}=1]$
$=\mathds{1}\{d_{t}=0,y_{t}=1\}\mathsf{E}[Y_{\tau+1}\mid D_{t}=0,Y_{t}=C_{t}=1,H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}}]$
$=r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t}).$
For $t<\tau$, we have
$\mathsf{m}_{t}(e_{t}^{\mathbbm{h}},v_{t})$ $=\mathsf{E}[\mathsf{m}_{t+1}(E_{t+1}^{\mathbbm{h}},V_{t+1})\mid D_{t}=d_{t},Y_{t}=y_{t},H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}},C_{t}=1]$
$=\mathsf{E}[R_{t+1}\mathsf{q}_{t+1}(A_{t+1}^{\mathbbm{d}},H_{t+1})\mid D_{t}=d_{t},Y_{t}=y_{t},H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}},C_{t}=1]$
$=r_{t}\mathsf{E}[R_{t+1}\mathsf{q}_{t+1}(A_{t+1}^{\mathbbm{d}},H_{t+1})\mid R_{t}=C_{t}=1,H_{t}=h_{t},A_{t}=a_{t}^{\mathbbm{d}}]$
$=r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t}).$
Since $R_{1}=1$, we have $\mathsf{m}_{1}(E_{1}^{\mathbbm{h}},V_{1})=\mathsf{q}_{1}(A_{1}^{\mathbbm{d}},H_{1})$, concluding the proof of the result. ∎
Appendix C Efficient influence function
Proof As before, let $V_{t}=(D_{t},Y_{t},L_{t})$ and $E_{t}=(A_{t},C_{t})$. Let $\mathbbm{h}(E_{t},V_{t})=(\mathbbm{d}(A_{t},H_{t}),1)$. First, notice that
$\mathsf{z}_{t}(e_{t},v_{t})$ $=\frac{\mathsf{g}_{t}^{\mathbbm{h}}(e_{t}\mid v_{t})}{\mathsf{g}_{t}(e_{t}\mid v_{t})}$
$=\frac{\mathsf{g}_{A,t}^{\mathbbm{d}}(a_{t}\mid h_{t})}{\mathsf{g}_{A,t}(a_{t}\mid h_{t})}\frac{c_{t}}{\mathsf{g}_{C,t}(a_{t},h_{t})}.$
Then we have
$\psi_{t}(z)$ $=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})\right)\{\mathsf{m}_{s+1}(e_{s+1}^{\mathbbm{h}},v_{s+1})-\mathsf{m}_{s}(e_{s},v_{s})\}+\mathsf{m}_{t}(e_{t}^{\mathbbm{h}},v_{t})$
$=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})-r_{s}\mathsf{q}_{s}(a_{s},h_{s})\}+r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})$
$=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})\right)r_{s}\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})-\mathsf{q}_{s}(a_{s},h_{s})\}+r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})$
$=\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{z}_{k}(e_{k},v_{k})r_{k}\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})-\mathsf{q}_{s}(a_{s},h_{s})\}+r_{t}\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})$
$=r_{t}\left[\sum_{s=t}^{\tau}\left(\prod_{k=t}^{s}\mathsf{w}_{t}(c_{k},a_{k},h_{k})\right)\{r_{s+1}\mathsf{q}_{s+1}(a_{s+1}^{\mathbbm{h}},h_{s+1})-\mathsf{q}_{s}(a_{s},h_{s})\}+\mathsf{q}_{t}(a_{t}^{\mathbbm{d}},h_{t})\right]$
$=r_{t}\varphi_{t}(x).$ (10)
An application of Theorem 3 together with the above results for $t=1$ yields the desired result after noticing that $R_{1}=1$. ∎
Appendix D von-Mises expansion
Lemma 1 (von-Mises expansion).
define
$\mathsf{H}_{t,s}^{\prime}=\prod_{l=t+1}^{s-1}\frac{\mathsf{g}_{A,t}^{\mathbbm{d}^{\prime}}(A_{t}\mid H_{t})}{\mathsf{g}_{A,t}^{\prime}(A_{t}\mid H_{t})}\frac{C_{t}}{\mathsf{g}_{C,t}^{\prime}(A_{t},H_{t})},$
Let
$\mathsf{J}_{t}(a_{t},h_{t})=\sum_{s=t+1}^{\tau}\mathsf{E}\left[\mathsf{H}_{t,s}^{\prime}R_{s}\{\mathsf{w}_{s}^{\prime}(A_{s},H_{s})-\mathsf{w}_{s}(A_{s},H_{s})\}\{\mathsf{q}_{s}^{\prime}(A_{s},H_{s})-\mathsf{q}_{s}(A_{s},H_{s})\}\mid R_{t}=C_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\right].$
Then, in the event $R_{t}=1$:
$\mathsf{q}_{t}(a_{t},h_{t})=\mathsf{E}[R_{t+1}\varphi_{t+1}^{\prime}(X)\mid R_{t}=C_{t}=1,A_{t}=a_{t},H_{t}=h_{t}]+\mathsf{J}_{t}(a_{t},h_{t}).$
Proof This results follows from an application of Theorem 4 to a data structure $V_{t}=(D_{t},Y_{t},L_{t})$ and $E_{t}=(A_{t},C_{t})$ with $\mathbbm{h}(E_{t},V_{t})=(\mathbbm{d}(A_{t},H_{t}),1)$, together with expression (10). First, note that
$\mathsf{U}_{t}(e_{t},v_{t})$ $=\sum_{s=t+1}^{\tau}\mathsf{E}\left[\mathsf{K}_{t,s}^{\prime}\{\mathsf{z}_{s}^{\prime}(E_{s},V_{s})-\mathsf{z}_{s}(E_{s},V_{s})\}\{\mathsf{m}_{s}^{\prime}(E_{s},V_{s})-\mathsf{m}_{s}(E_{s},V_{s})\}\mid E_{t}=e_{t},V_{t}=v_{t}\right]$
$=\sum_{s=t+1}^{\tau}\mathsf{E}\left[\mathsf{H}_{t,s}^{\prime}R_{s}\{\mathsf{w}_{s}^{\prime}(A_{s},H_{s})-\mathsf{w}_{s}(A_{s},H_{s})\}\{\mathsf{q}_{s}^{\prime}(A_{s},H_{s})-\mathsf{q}_{s}(A_{s},H_{s})\}\mid D_{t}=d_{t},Y_{t}=y_{t},C_{t}=c_{t},A_{t}=a_{t},H_{t}=h_{t}\right]$
$=\sum_{s=t+1}^{\tau}\mathsf{E}\left[\mathsf{H}_{t,s}^{\prime}R_{s}\{\mathsf{w}_{s}^{\prime}(A_{s},H_{s})-\mathsf{w}_{s}(A_{s},H_{s})\}\{\mathsf{q}_{s}^{\prime}(A_{s},H_{s})-\mathsf{q}_{s}(A_{s},H_{s})\}\mid D_{t}=d_{t},Y_{t}=y_{t},C_{t}=c_{t},A_{t}=a_{t},H_{t}=h_{t}\right]$
$=r_{t}\sum_{s=t+1}^{\tau}\mathsf{E}\left[\mathsf{H}_{t,s}^{\prime}R_{s}\{\mathsf{w}_{s}^{\prime}(A_{s},H_{s})-\mathsf{w}_{s}(A_{s},H_{s})\}\{\mathsf{q}_{s}^{\prime}(A_{s},H_{s})-\mathsf{q}_{s}(A_{s},H_{s})\}\mid R_{t}=C_{t}=1,A_{t}=a_{t},H_{t}=h_{t}\right]$
$=r_{t}\mathsf{J}_{t}(a_{t},h_{t})$
Specifically, we have
$r_{t}\mathsf{q}_{t}(a_{t},h_{t})$ $=\mathsf{m}_{t}(e_{t},h_{t})$
$=\mathsf{E}[\psi_{t+1}^{\prime}(Z)\mid E_{t}=e_{t},V_{t}=v_{t}]+\mathsf{U}_{t}(e_{t},v_{t})$
$=\mathsf{E}[R_{t+1}\varphi_{t+1}^{\prime}(X)\mid D_{t}=d_{t},Y_{t}=y_{t},C_{t}=c_{t},A_{t}=a_{t},H_{t}=h_{t}]+\mathsf{U}_{t}(e_{t},v_{t})$
$=r_{t}\mathsf{E}[R_{t+1}\varphi_{t+1}^{\prime}(X)\mid R_{t}=C_{t}=1,A_{t}=a_{t},H_{t}=h_{t}]+r_{t}\mathsf{J}_{t}(a_{t},h_{t}),$
which concludes the proof. ∎
Appendix E Targeted Minimum Loss Based Estimator
The targeted minimum loss-based estimator is computed as a substitution estimator that uses an estimate $\tilde{\mathsf{q}}_{1,j(i)}$ carefully constructed to solve the cross-validated efficient influence function estimating equation $\mathsf{P}_{n}\{\varphi_{1}(\cdot,\tilde{\eta}_{j(\cdot)})-\hat{\theta}_{\mbox{tmle}}\}=0$. The construction of $\tilde{\mathsf{q}}_{1,j(i)}$ is motivated by the observation that the efficient influence function of $\theta$ can be expressed as a sum of terms of the form:
$\left(\prod_{k=1}^{t}\mathsf{w}_{k}(c_{k},a_{k},h_{k})\right)\{r_{t+1}\mathsf{q}_{t+1}(a_{t+1}^{\mathbbm{d}},h_{t+1})-\mathsf{q}_{t}(a_{t},h_{t})\},$
which take the form of score functions $\lambda(W)\{M-\mathsf{E}(M\mid W)\}$ for appropriately defined variables $M$ and $W$ and some weight function $\lambda$. It is well known that if $\mathsf{E}(M\mid W)$ is estimated within a weighted generalized linear model with canonical link that includes an intercept, then the weighted MLE estimate solves the score equation $\sum_{i}\lambda(W_{i})\{M_{i}-\hat{\mathsf{E}}(M_{i}\mid W_{i})\}=0$. TMLE uses this observation to iteratively tilt preliminary estimates of $\mathsf{q}_{t}$ towards a solution of the efficient influence function estimating equation. The algorithm is defined as follows:
  1. Step 1.
    Initialize $\tilde{\eta}=\hat{\eta}$ and $\tilde{\mathsf{q}}_{\tau+1,j(i)}(A^{\mathbbm{d}}_{\tau+1,i},H_{\tau+1,i})=Y_{i}$.
  2. Step 2.
    For $s=1,\ldots,\tau$, compute the weights
    $\lambda_{s,i}=\prod_{k=1}^{s}\hat{\mathsf{w}}_{k,j(i)}(C_{k,i},A_{k,i},H_{k,i})$
  3. Step 3.
    For $t=\tau,\ldots,1$:
    • Compute the pseudo-outcome $\check{Y}_{t+1,i}=R_{t+1}\tilde{\mathsf{q}}_{t+1,j(i)}(A^{\mathbbm{d}}_{t+1,i},H_{t+1,i})$
    • Fit the generalized linear tilting model
      $\operatorname{link}\tilde{\mathsf{q}}_{t}^{\epsilon}(A_{t,i},H_{t,i})=\epsilon+\operatorname{link}\tilde{\mathsf{q}}_{t,j(i)}(A_{t,1},H_{t,i})$
      where $\operatorname{link}(\cdot)$ is the canonical link. Here, $\operatorname{link}\tilde{\mathsf{q}}_{t,j(i)}(a_{t},h_{t})$ is an offset variable (i.e., a variable with known parameter value equal to one). The parameter $\epsilon$ may be estimated by running a generalized linear model of the pseudo-outcome $\check{Y}_{t+1,i}$ with only an intercept term, an offset term equal to $\operatorname{link}\tilde{\mathsf{q}}_{t,j(i)}(A_{t,i},H_{t,i})$, and weights $\lambda_{t,i}$, using all the data points in the sample such that $R_{i,t}=C_{i,t}=1$. An outcome bounded in an interval $[a,b]$ may be analyzed with logistic regression (i.e., $\operatorname{link}=\operatorname{logit}$) by mapping it to an outcome $(0,1)$ through the transformation $(Y-a)/(b-a)(1-2\gamma)+\gamma$ for some small value $\gamma>0$. This approach has robustness advantages compared to fitting a linear model as it guarantees that the predictions in the next step remain within the outcome space (see Gruber & van der Laan, 2010).
    • Let $\hat{\epsilon}$ denote the maximum likelihood estimate, and update the estimates as
      $\operatorname{link}\tilde{\mathsf{q}}_{t,j(i)}^{\hat{\epsilon}}(A_{t,i},H_{t,i})$ $=\hat{\epsilon}+\operatorname{link}\tilde{\mathsf{q}}_{t,j(i)}(A_{t,i},H_{t,i})$
      $\operatorname{link}\tilde{\mathsf{q}}_{t,j(i)}^{\hat{\epsilon}}(A_{t,i}^{\mathbbm{d}},H_{t,i})$ $=\hat{\epsilon}+\operatorname{link}\tilde{\mathsf{q}}_{t,j(i)}(A_{t,i}^{\mathbbm{d}},H_{t,i}).$
      The above procedure with canonical link guarantees the following score equation is solved:
      $\frac{1}{n}\sum_{i=1}^{n}\left(\prod_{k=1}^{t}\hat{\mathsf{w}}_{k,j(i)}(C_{k,i},A_{k,i},H_{k,i})\right)\{R_{t+1,i}\tilde{\mathsf{q}}_{t+1,j(i)}(A^{\mathbbm{d}}_{t+1,i},H_{t+1,i})-\tilde{\mathsf{q}}_{t,j(i)}^{\hat{\epsilon}}(A_{t,i},H_{t,i})\}=0$
    • Update $\tilde{\mathsf{q}}_{t,j(i)}=\tilde{\mathsf{q}}_{t,j(i)}^{\hat{\epsilon}}$, $t=t-1$, and iterate.
  4. Step 4.
    The TMLE is defined as
    $\hat{\theta}=\frac{1}{n}\sum_{i=1}^{n}\tilde{\mathsf{q}}_{1,j(i)}(A_{1,i}^{\mathbbm{d}},L_{1,i}).$
The iterative procedure and the score equation argument above guarantee that
$\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{\tau}\left(\prod_{k=1}^{t}\hat{\mathsf{w}}_{k,j(i)}(C_{k,i},A_{k,i},H_{k,i})\right)\{R_{t+1,i}\tilde{\mathsf{q}}_{t+1,j(i)}(A_{t+1,i}^{\mathbbm{d}},H_{t+1,i})-\tilde{\mathsf{q}}_{t,j(i)}(A_{t,i},H_{t,i})\}=0,$
and thus that $\mathsf{P}_{n}\{\varphi_{1}(\cdot;\tilde{\eta}_{j(\cdot)})-\hat{\theta}\}=0$. This fact is crucial to prove weak convergence and consistency of the TMLE.
Arguing along the lines of the proofs presented in Díaz et al. (2021), $\sqrt{n}$-consistency of the above TMLE will require that the weights $\mathsf{w}_{t}$ as well as their estimators are all bounded above, and that the second order term
$\sum_{t=1}^{\tau}\lVert\hat{\mathsf{w}}_{t}-\mathsf{w}_{t}\rVert\lVert\tilde{\mathsf{q}}_{t}-\mathsf{q}_{t}\rVert$ (11)
is $o_{P}(n^{-1/2})$. Note that, for each $t$, we have
$\lVert\tilde{\mathsf{q}}_{t-1}-\mathsf{q}_{t-1}\rVert$ $\leq\lVert\tilde{\mathsf{q}}_{t-1}-\mathsf{q}_{t-1}^{\star}\rVert+\lVert\tilde{\mathsf{q}}_{t-1}^{\star}-\mathsf{q}_{t-1}\rVert$
$=\lVert\tilde{\mathsf{q}}_{t-1}-\mathsf{q}_{t-1}^{\star}\rVert+O_{P}(\lVert\tilde{\mathsf{q}}_{t}-\mathsf{q}_{t}\rVert).$
Recursive application of the above for $t=\tau-1,\ldots,2$ means that the term (11) is bounded in probability by
$\sum_{t=1}^{\tau}\left(\lVert\hat{\mathsf{w}}_{t}-\mathsf{w}_{t}\rVert\sum_{s=t}^{\tau}\lVert\hat{\mathsf{q}}_{s}-\mathsf{q}_{s}^{\star}\rVert\right).$
Convergence of this term at rate $o_{P}(n^{-1/2})$ seems to require convergence of cross-product terms of the type $\lVert\hat{\mathsf{w}}_{t}-\mathsf{w}_{t}\rVert\lVert\hat{\mathsf{q}}_{s}-\mathsf{q}_{s}^{\star}\rVert$ for $s\geq t$, which are not required by the SDR estimator (see Theorem 1 in the main document).
Appendix F Distribution of censoring weights in the application
Figure 3: Distribution of the estimated cumulative censoring probabilities $\prod_{s=1}^{t}\hat{\mathsf{g}}_{C}(A_{t,i},H_{t-i})$ for each time $t$.